To TV stations offer deals for putting ads on. Station 1’s rate is $20 intail charge, plus $1 per day and stations 2’s rate is $8 intail charge, plus $2 per day. On what day will both options equal each other in value. Create a system of equations and use the equal value theorem to solve

Answer by math_tutor2020(3835)


x = number of days
y = amount charged

Station 1:
y = x+20

Station 2:
y = 2x+8

Apply substitution to solve
2x+8 = x+20
2x-x = 20-8
x = 12
Both stations charge the same amount if 12 days are used.

Confirmation:
y = x+20 = 12+20 = 32
y = 2x+8 = 2*12+8 = 24+8 = 32
Both stations charge $32, so the answer is confirmed.